OpenStudy (anonymous):

When 135 grams of titanium tetrachloride react with an excess of water, as shown in the unbalanced chemical equation below, how many grams of hydrochloric acid will be produced? TiCl4 + H2O TiO2 + HCl

OpenStudy (anonymous):

First you need to balance the equation: TiCl4 +H2O = TiO2 + HCL (unbalanced) TiCl4 +2H2O = TiO2 + 4HCL (balanced)

OpenStudy (anonymous):

okay than what?

OpenStudy (anonymous):

Then you need to take what's given to you in the problem and convert that to mole in this case it's 135 grams of TiCl4

OpenStudy (anonymous):

then you need to figure out the ratio between the mol of TiCl4 and HCL, this case it should be from the balanced equation which is 4:1 ratio

OpenStudy (anonymous):

final step: you need to convert from mole to grams cuz of what the question asking ( in grams) convert from mol HCL to grams of HCL

OpenStudy (anonymous):

let me write out the whole step for you hold on..

OpenStudy (anonymous):

Start with 135g TiCl4 ( 1mol TiCl4/ 189.7 g TiCl4)(4 mol HCL/1 mol HCL)(36.45g HCL/1mol HCl) = 103.7 g HCL ~ 104 g HCL

OpenStudy (anonymous):

if you write out what i did on paper, it would be easier for you to see what's happening...

OpenStudy (anonymous):

the 189.7 came from the TiCl4 molar mass how to find molar mass: Ti = 47.90 and Cl4 = 35.45x4 =141.8 then juts add the two 47.90 + 141.8 = 189.7 g TiCl4

OpenStudy (anonymous):

and the 36.45 came from molar mass of HCl

OpenStudy (anonymous):

okay so the final answer is 104g of HCI right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright thank you so much(:

OpenStudy (anonymous):

np :D